# Optimal navigation in two-dimensional regular and turbulent flows

### v1

@@ Abstract
Zermelo’s navigation problem seeks the trajectory of minimal travel time between two points in a fluid flow. We address this problem for an agent – such as a micro-robot or active particle – that is advected by a two-dimensional flow, self-propels at a fixed speed smaller than or comparable to the characteristic flow velocity, and can steer its direction. The flows considered span increasing levels of complexity, from steady solid-body rotation to the Taylor-Green flow and fully developed turbulence in the inverse cascade regime. Although optimal control theory provides time-minimizing trajectories, these solutions become unstable in chaotic regimes realized for complex background flows. To design robust navigation strategies under such conditions, we apply reinforcement learning. Both action-value (Q-learning) and policy-gradient (one-step actor-critic) methods achieve successful navigation with comparable performance. Crucially, we show that agents trained on coarse-grained flows – retaining only large-scale features – generalize effectively to the full turbulent field. This robustness to incomplete flow information is essential for practical navigation in real-world oceanic and atmospheric environments.
@@ I Introduction
The study of Zermelo’s problem has traditionally been approached through the lens of optimal control theory. Within this framework, one can derive a differential equation governing the evolution of the agent’s steering direction, which serves as the necessary condition for optimality (see, e.g., Ref. [bryson2018applied, Sec. 2.7]). For relatively simple flow fields, this equation admits analytical solutions [techy2011optimal], enabling the construction of optimal trajectories that can be used as benchmarks. In more complex environments, however, numerical techniques are required. The challenge is further complicated by the absence of a known initial condition for the steering equation, which in practice must be determined using shooting methods [kirk2004optimal].
@@
This classical formulation faces two crucial limitations. First, calculating the optimal path requires perfect and global knowledge of the flow field. Second, the resulting optimal trajectories can be unstable. In systems exhibiting chaotic dynamics – characterized by exponential separation of nearby trajectories – infinitesimal uncertainties in initial conditions or control inputs are amplified, leading to drastic deviations from the planned route [angelo2009chaos, biferale2019zermelo]. In these cases, pre-computed optimal control strategies become ineffective, as the agent must perform additional actions to compensate deviations from the found path. Both limitations are violated in turbulent flows, which dominate oceanic and atmospheric environments [vallis2017atmospheric, boffetta2012two]. The requirement for perfect flow knowledge is unrealistic, as operational forecasts are inherently coarse-grained, forcing an agent to operate with incomplete information. Furthermore, turbulence is a canonical example of a chaotic system, meaning the stability condition is violated.
@@
In response to these challenges, the field has explored alternative paradigms, including heuristic [piro2022optimal, monthiller2022surfing, calascibetta2023optimal] and data-driven methods [yoo2016path, colabrese2017flow, gustavsson2017finding, schneider2019optimal, biferale2019zermelo, buzzicotti2020optimal, alageshan2020machine, gunnarson2021learning, nasiri2023optimal]. The rise of machine learning has opened new avenues, with reinforcement learning (RL) emerging as a particularly powerful model-free framework for complex control [sutton1998reinforcement]. The RL formulation treats navigation as a sequential decision-making process in which an agent learns a policy – a mapping from states (e.g., position) to actions (e.g., heading adjustments) – by maximizing a cumulative reward, typically defined in terms of minimizing travel time to a target. Early studies have highlighted the promise of this approach. For instance, the work of Colabrese et al. [colabrese2017flow] demonstrated that smart active particles could use Q-learning to navigate vortex flows, effectively discovering physical strategies like surfing on specific flow structures. Similarly, Biferale et al. [biferale2019zermelo] explored the use of RL to solve Zermelo’s problem in synthetic two-dimensional turbulent flows, showing that agents can learn to leverage flow structures to minimize travel time. However, a systematic investigation across a hierarchy of flows from regular to fully turbulent, and crucially, an assessment of the robustness and generalization capabilities of learned policies under realistic information constraints, remains an open area of research.
@@
By maintaining a detailed presentation, our work serves a dual purpose: advancing the research frontier and providing a valuable reference for educational purposes. The code that supports the findings of this study are openly available at https://github.com/parfenyev/opt-navig. The remainder of this paper is organized as follows. In Section LABEL:sec:2, we formulate the Zermelo’s problem and introduce the equations of motion for the agent. Section LABEL:sec:3 describe the hierarchy of flows we study. Section LABEL:sec:4 details the application of optimal control theory and analyzes the stability of the resulting trajectories. Section LABEL:sec:5 explains our reinforcement learning framework, including the state and action spaces, reward structure, and the implementation details of the Q-learning and one-step actor-critic algorithms. Section LABEL:sec:6 is devoted to navigation in turbulent environment with incomplete flow information. Finally, Section LABEL:sec:7 summarizes our findings.
@@ II Problem Setup
[…]
@@ III.2 Steady sink and time-varying vortex flow
[…]
The flow model represents a constant sink plus a vortex that linearly changes vorticity with time. The goal is to start from the point $(x_{A},y_{A})=\left(1/2,\;\sqrt{3}/2\right)$ and reach the point $(x_{B},y_{B})=(1,\;0)$ in the shortest time. The agent’s speed is $V_{0}=1$.
@@ III.3 Taylor-Green vortices
$label{eq:TG-flow}\bm{v}(x,y)=\begin{pmatrix}u_{0}\sin(kx)\cos(ky)\\ -u_{0}\cos(kx)\sin(ky)\end{pmatrix}$. (4)
The corresponding vorticity is $\varpi=\partial_{x}v_{y}-\partial_{y}v_{x}=2u_{0}k\sin(kx)\sin(ky)$. Thus, the velocity field looks like a checkerboard of counter-rotating vortices. A similar velocity field can be generated on the fluid surface due to the nonlinear interaction of two orthogonal standing waves [filatov2016nonlinear, filatov2016generation, francois2017wave, xia2019generation, abella2020measurement].
@@
$label{eq:NS-equation}\partial_{t}\bm{v}+(\bm{v}\cdot\nabla)\bm{v}=-\nabla p+\nu\nabla^{2}\bm{v}-\alpha\bm{v}+\bm{f},\quad\nabla\cdot\bm{v}=0,$ (5)
@@
DNS results are obtained by integrating (LABEL:eq:NS-equation) using the GeophysicalFlows.jl pseudospectral code [GeophysicalFlowsJOSS], fully dealiazed by the two-thirds rule. We use a $512^{2}$ grid with a fixed integration time step $dt=2\cdot 10^{-4}$, resulting in a Courant number <0.3<0.3. The forcing correlation time is also $dt$. Starting from rest, the simulation is run until it reaches a non-equilibrium stationary state, observed by the saturation of the total kinetic energy. Snapshots of the velocity field are then output at an interval of $\tau=5\cdot 10^{-3}$ over a total duration of $T_{tot}=6.15$, yielding $1230$ snapshots for analysis.
The resulting flow is characterized by dimensionless numbers $Re=UL/\nu\approx 8\cdot 10^{3}$ and $Rh=U/(\alpha L)\approx 5$, where $U\approx 6.3$ is the root-mean-square velocity. The mean energy spectrum, shown in Fig. LABEL:fig:1, confirms the established phenomenology of forced 2D turbulence. In the inverse energy cascade range $k<k_{f}$, the spectrum follows the theoretical $E(k)\propto k^{-5/3}$ scaling [boffetta2012two], while in the direct enstrophy cascade range $k>k_{f}$, the bottom drag leads to a steeper slope than the classical $E(k)\propto k^{-3}$ prediction [valadao2025spectrum].
@@ IV Optimal Control Theory
IV.1 Zermelo equation
The problem of finding the minimum-time trajectory and the corresponding travel time $T$ can be addressed through variational calculus. We seek to minimize the functional $\int_{0}^{T}dt\;L(t,\bm{R}_{t},\dot{\bm{R}}_{t})$, where $\bm{R}_{t}$ obeys the equation of motion (LABEL:eq:motion) and boundary conditions $\bm{R}_{0}=\bm{r}_{A}$ and $\bm{R}_{T}=\bm{r}_{B}$. This constitutes a constrained optimization problem that can be solved using the method of Lagrange multipliers. For the minimum-time problem, the cost function is simply $L(t,\bm{R}_{t},\dot{\bm{R}}_{t})\equiv 1$, yielding the extended functional
@@ IV.1 Zermelo equation
The obtained system is governed by four differential equations (for $\bm{R}_{t}$ and $\bm{\lambda}_{t}$) with four boundary conditions (the fixed start and end points, $\bm{R}_{0}=\bm{r}_{A}$ and $\bm{R}_{T}=\bm{r}_{B}$), one optimality condition for the control $\theta_{t}$, and one transversality condition $\bm{\lambda}_{T}\cdot\dot{\bm{R}}_{T}=1$ for the free terminal time $T$. The state equation $\dot{\bm{R}}_{t}=\partial H/\partial\bm{\lambda}_{t}$ simply restates the original equations of motion (LABEL:eq:motion). From $\dot{\bm{\lambda}}_{t}=-\partial H/\partial\bm{R}_{t}$ and $\partial H/\partial\theta_{t}=0$, we obtain
@@
The optimal control is determined by $\theta_{t}=\arctan(\lambda_{t}^{y}/\lambda_{t}^{x})$. A more useful form is the Zermelo equation, derived by differentiating this expression with respect to time
[…]
@@
The complete Two-Point Boundary Value Problem (TPBVP) is thus defined by coupling the equations of motion (LABEL:eq:motion) with the heading evolution equation (LABEL:eq:angle). The boundary conditions are the fixed start and end positions, $\bm{R}_{0}=\bm{r}_{A}$ and $\bm{R}_{T}=\bm{r}_{B}$. The initial heading $\theta_{0}$ is not specified, making this a free condition. The standard numerical approach to solve this TPBVP is the shooting method: an initial guess for $\theta_{0}$ is iteratively refined until the resulting trajectory, obtained by forward integration of the system, satisfies the terminal condition $\bm{R}_{T}=\bm{r}_{B}$ within a desired tolerance. More details on the numerical procedure will be provided later when discussing specific examples. Note that once the initial heading $\theta_{0}$ is known, the optimal path can be integrated autonomously, provided the agent can measure the local velocity field gradients, see Eq. (LABEL:eq:angle).
@@ IV.2 Example 1: Steady vortex flow
[…]
sec:OPC_A
For the background velocity field $\bm{v}(x,y)=\left(-\omega y,\;\omega x\right)^{T}$, the Zermelo equation (LABEL:eq:angle) has a simple solution $\theta_{t}=\theta_{0}+\omega t$. Integrating the equations of motion (LABEL:eq:motion) from the starting point $(X_{0},Y_{0})=(x_{A},y_{A})$ yields the trajectory:
@@
The initial heading $\theta_{0}$ and total travel time $T$ are determined by enforcing the terminal condition $(X_{T},Y_{T})=(x_{B},y_{B})$. For the numerical parameters specified in Section LABEL:sec:Ex_A, we obtain:
@@
The found optimal trajectory is illustrated in Fig. LABEL:fig:2 by a white line. The figure also includes trajectories corresponding to small deviations in the initial heading $\theta_{0}$ and slight perturbations in the initial position $\bm{r}_{A}$. Overall, these trajectories stay close to the optimal path and successfully reach a narrow neighborhood of the target point. Based on expressions (IV.2) and (IV.2), it can be concluded that the initial perturbations exhibit linear growth with time rather than exponential growth, indicating the stability of the optimal trajectory.
[…]
@@ IV.3 Example 2: Steady sink and time-varying vortex flow
[…]
sec:OPC_B
For the background velocity field (LABEL:eq:sink+rotation), the Zermelo equation (LABEL:eq:angle) has a solution $\theta_{t}=\theta_{0}+t/2-t^{2}/2$. Then, the equations of motion (LABEL:eq:motion) can be integrated from the starting point $(X_{0},Y_{0})=(x_{A},y_{A})$
@@
The found trajectory should also pass through the terminal point $(X_{T},Y_{T})=(x_{B},y_{B})$, which allows us to obtain the initial heading $\theta_{0}$ and travel time $T$. For the numerical parameters specified in Section LABEL:sec:Ex_B, we obtain:
@@
Figure LABEL:fig:3 presents the optimal trajectory (white line) superimposed on the velocity field at the final time $T$. The figure also illustrate the system’s sensitivity by displaying trajectories with small errors in the initial heading $\theta_{0}$ and initial position $\bm{r}_{A}$. All trajectories remain near the optimal path and successfully approach the destination point. The complete time-dependent evolution is provided in the video, see Ref. [repository]. Based on expressions (18) and (19), it can be concluded that the initial small perturbations $\delta X_{0},\delta Y_{0},\delta\theta_{0}$ do not exhibit exponential growth with time, indicating the stability of the optimal trajectory.
[…]

### v2

# Optimal navigation in two-dimensional flows: control theory and reinforcement learning

@@ Abstract
Zermelo’s navigation problem seeks the trajectory of minimal travel time between two points in a fluid flow. We address this problem for an agent – such as a micro-robot or active particle – that is advected by a two-dimensional flow, self-propels at a fixed speed smaller than or comparable to the characteristic flow velocity, and can steer its direction. The flows considered span increasing levels of complexity, from steady solid-body rotation to the Taylor-Green flow and fully developed turbulence in the inverse cascade regime. Although optimal control theory provides time-minimizing trajectories, these solutions become unstable in chaotic regimes realized for complex background flows. To design robust navigation strategies under such conditions, we apply reinforcement learning. Both action-value (Q-learning) and policy-gradient (one-step actor-critic) methods achieve successful navigation with comparable performance. Crucially, we show that agents trained on coarse-grained turbulent flows – retaining only large-scale features – generalize effectively to the full velocity field. This robustness to incomplete flow information is essential for practical navigation in real-world oceanic and atmospheric environments.
@@ I Introduction
The study of Zermelo’s problem has traditionally been approached through the lens of optimal control theory. Within this framework, one can derive a differential equation governing the evolution of the agent’s steering direction, which serves as the necessary condition for optimality (see, e.g., Ref. [bryson2018applied, Sec. 2.7]). For relatively simple flow fields, this equation admits analytical solutions [techy2011optimal], enabling the construction of optimal trajectories. In more complex environments, however, numerical techniques are required. The challenge is further complicated by the absence of a known initial condition for the steering equation, which in practice must be determined using shooting methods [kirk2004optimal].
@@
This classical formulation faces two crucial limitations. First, calculating the optimal path requires perfect and global knowledge of the flow field. Second, the resulting optimal trajectories can be unstable. In systems exhibiting chaotic dynamics – characterized by exponential separation of nearby trajectories – infinitesimal uncertainties in initial conditions or control inputs are amplified, leading to drastic deviations from the planned route [angelo2009chaos, biferale2019zermelo]. In these cases, pre-computed optimal control strategies become ineffective. Both limitations are violated in turbulent flows, which dominate oceanic and atmospheric environments [vallis2017atmospheric, boffetta2012two]. The requirement for perfect flow knowledge is unrealistic, as operational forecasts are inherently coarse-grained, forcing an agent to operate with incomplete information. Furthermore, turbulence is a canonical example of a chaotic system, meaning the stability condition is violated.
@@
In response to these challenges, the field has explored alternative paradigms, including heuristic [piro2022optimal, monthiller2022surfing, calascibetta2023optimal] and data-driven methods [yoo2016path, colabrese2017flow, gustavsson2017finding, schneider2019optimal, biferale2019zermelo, buzzicotti2020optimal, alageshan2020machine, gunnarson2021learning, nasiri2023optimal]. The rise of machine learning has opened new avenues, with reinforcement learning (RL) emerging as a particularly powerful model-free framework for complex control [sutton1998reinforcement]. The RL formulation treats navigation as a sequential decision-making process in which an agent learns a policy – a mapping from states (e.g., positions) to actions (e.g., heading adjustments) – by maximizing a cumulative reward, typically defined in terms of minimizing travel time to a target. Early studies have highlighted the promise of this approach. For instance, the work of Colabrese et al. [colabrese2017flow] demonstrated that smart active particles could use Q-learning to navigate vortex flows, effectively discovering physical strategies like surfing on specific flow structures. Similarly, Biferale et al. [biferale2019zermelo] explored the use of RL to solve Zermelo’s problem in synthetic two-dimensional turbulent flows, showing that agents can learn to leverage flow structures to minimize travel time. However, a systematic investigation across a hierarchy of flows from regular to fully turbulent, and crucially, an assessment of the robustness and generalization capabilities of learned policies under realistic information constraints, remains an open area of research.
@@
By maintaining a detailed presentation, our work serves a dual purpose: advancing the research frontier and providing a valuable reference for educational purposes. The code that supports the findings of this study are openly available in Ref. [repository]. The remainder of this paper is organized as follows. In Section II, we formulate the Zermelo’s problem and introduce the equations of motion for the agent. Section III describe the hierarchy of flows we study. Section IV details the application of optimal control theory and analyzes the stability of the resulting trajectories. Section V explains our reinforcement learning framework and discusses its application to the presented examples. Section VI is devoted to navigation in turbulent environment with incomplete flow information. Finally, Section VII summarizes our findings.
@@ II Problem Setup
$\dfrac{d\bm{R}_{t}}{dt}=\bm{v}(\bm{R}_{t},t)+V_{0}\begin{pmatrix}\cos\theta_{t}\\ \sin\theta_{t}\end{pmatrix}$, (1)
@@ III.2 Steady sink and time-varying vortex flow
$\bm{v}(x,y,t)=\begin{pmatrix}-0.3x-(0.5-t)y\\ -0.3y+(0.5-t)x\end{pmatrix}$. (3)
The flow model represents a constant sink plus a vortex that linearly changes vorticity with time. The goal is to start from the point $(x_{A},y_{A})=\left(1/2,\;\sqrt{3}/2\right)$ and reach the point $(x_{B},y_{B})=(1,\;0)$ in the shortest time. The agent’s speed is $V_{0}=1$. The main purpose of this example is to demonstrate the applicability of the discussed approaches to time-varying and divergent flows.
@@ III.3 Taylor-Green vortices
$\bm{v}(x,y)=\begin{pmatrix}u_{0}\sin(kx)\cos(ky)\\ -u_{0}\cos(kx)\sin(ky)\end{pmatrix}$. (4)
The corresponding vorticity is $\varpi=\partial_{x}v_{y}-\partial_{y}v_{x}=2u_{0}k\sin(kx)\sin(ky)$. Thus, the velocity field looks like a checkerboard of counter-rotating vortices. A similar velocity field can be generated on the fluid surface due to the nonlinear interaction of two orthogonal standing waves [filatov2016nonlinear, filatov2016generation, francois2017wave, xia2019generation, parfenyev2019formation, abella2020measurement].
@@
$\partial_{t}\bm{v}+(\bm{v}\cdot\nabla)\bm{v}=-\nabla p+\nu\nabla^{2}\bm{v}-\alpha\bm{v}+\bm{f},\quad\nabla\cdot\bm{v}=0,$ (5)
@@
DNS results are obtained by integrating (5) using the GeophysicalFlows.jl pseudospectral code [GeophysicalFlowsJOSS], fully dealiazed by the two-thirds rule. We use a $512^{2}$ grid with a fixed integration time step $dt=2\cdot 10^{-4}$, resulting in a Courant number <0.3<0.3. The forcing correlation time is also $dt$. Starting from rest, the simulation is run until it reaches a non-equilibrium stationary state, observed by the saturation of the total kinetic energy. Snapshots of the velocity field are then output at an interval of $\tau=5\cdot 10^{-3}$ over a total duration of $T_{tot}=6.15$, yielding $1230$ snapshots for analysis.
The resulting flow is characterized by dimensionless numbers $Re=UL/\nu\approx 8\cdot 10^{3}$ and $Rh=U/(\alpha L)\approx 5$, where $U\approx 6.3$ is the root-mean-square velocity. The mean energy spectrum, shown in Fig. 1, confirms the established phenomenology of forced 2D turbulence. In the inverse energy cascade range $k<k_{f}$, the spectrum follows the theoretical $E(k)\propto k^{-5/3}$ scaling [boffetta2012two], while in the direct enstrophy cascade range $k>k_{f}$, the bottom drag leads to a steeper slope than the classical $E(k)\propto k^{-3}$ prediction [valadao2025spectrum].
@@ IV Optimal Control Theory
IV.1 Zermelo’s equation
The problem of finding the minimum-time trajectory and the corresponding travel time $T$ can be addressed through variational calculus. We seek to minimize the functional $\int_{0}^{T}dt\;L(t,\bm{R}_{t},\dot{\bm{R}}_{t})$, where $\bm{R}_{t}$ obeys the equation of motion (1) and boundary conditions $\bm{R}_{0}=\bm{r}_{A}$ and $\bm{R}_{T}=\bm{r}_{B}$. This constitutes a constrained optimization problem that can be solved using the method of Lagrange multipliers. For the minimum-time problem, the cost function is simply $L(t,\bm{R}_{t},\dot{\bm{R}}_{t})\equiv 1$, yielding the extended functional
@@ IV.1 Zermelo’s equation
The obtained system is governed by four differential equations (for $\bm{R}_{t}$ and $\bm{\lambda}_{t}$) with four boundary conditions (the fixed start and end points, $\bm{R}_{0}=\bm{r}_{A}$ and $\bm{R}_{T}=\bm{r}_{B}$), one optimality condition for the control $\theta_{t}$, and one transversality condition $\bm{\lambda}_{T}\cdot\dot{\bm{R}}_{T}=1$ for the free terminal time $T$. The state equation $\dot{\bm{R}}_{t}=\partial H/\partial\bm{\lambda}_{t}$ simply restates the original equations of motion (1). From $\dot{\bm{\lambda}}_{t}=-\partial H/\partial\bm{R}_{t}$ and $\partial H/\partial\theta_{t}=0$, we obtain
@@
The optimal control is determined by $\theta_{t}=\arctan(\lambda_{t}^{y}/\lambda_{t}^{x})$. A more useful form is the Zermelo’s equation, derived by differentiating this expression with respect to time and using expressions (11) and (12):
$\begin{split}\dot{\theta}_{t}=\partial_{x}v_{y}\sin^{2}\theta_{t}-\partial_{y}v_{x}\cos^{2}\theta_{t}\\ +(\partial_{x}v_{x}-\partial_{y}v_{y})\sin\theta_{t}\cos\theta_{t},\end{split}$ (14)
@@
The complete Two-Point Boundary Value Problem (TPBVP) is thus defined by coupling the equations of motion (1) with the heading evolution equation (14). The boundary conditions are the fixed start and end positions, $\bm{R}_{0}=\bm{r}_{A}$ and $\bm{R}_{T}=\bm{r}_{B}$. The initial heading $\theta_{0}$ is not specified, making this a free condition. The standard numerical approach to solve this TPBVP is the shooting method: an initial guess for $\theta_{0}$ is iteratively refined until the resulting trajectory, obtained by forward integration of the system, satisfies the terminal condition $\bm{R}_{T}=\bm{r}_{B}$ within a desired tolerance. More details on the numerical procedure will be provided later when discussing specific examples. Note that once the initial heading $\theta_{0}$ is known, the optimal path can be integrated autonomously, provided the agent can measure the local velocity field gradients, see Eq. (14).
@@ IV.2 Example 1: Steady vortex flow
For the background velocity field $\bm{v}(x,y)=\left(-\omega y,\;\omega x\right)^{T}$, the Zermelo’s equation (14) has a simple solution $\theta_{t}=\theta_{0}+\omega t$. Integrating the equations of motion (1) from the starting point $(X_{0},Y_{0})=(x_{A},y_{A})$ yields the trajectory:
@@
The initial heading $\theta_{0}$ and total travel time $T$ are determined by enforcing the terminal condition $(X_{T},Y_{T})=(x_{B},y_{B})$. For the numerical parameters specified in Section III.1, we obtain:
@@
The found optimal trajectory is illustrated in Fig. 2 by a white line. The figure also includes trajectories corresponding to small deviations in the initial heading $\theta_{0}$ and slight perturbations in the initial position $\bm{r}_{A}$. Overall, these trajectories stay close to the optimal path on the considered time scale and successfully reach a narrow neighborhood of the target point. Based on expressions (IV.2) and (IV.2), it can be concluded that the initial perturbations exhibit linear growth with time rather than exponential growth, indicating only the weak instability of the optimal trajectory.
Figure 2: Optimal control solution (white line) for the steady vortex flow, see Sections III.1. Near-optimal trajectories are shown in cyan and blue for the initial control angle $\theta_{0}\pm\pi/360$, and in magenta and red for the initial position error $\Delta=0.025$ in $x_{A}$ and $y_{A}$.
@@ IV.3 Example 2: Steady sink and time-varying vortex flow
For the background velocity field (3), the Zermelo’s equation (14) has a solution $\theta_{t}=\theta_{0}+t/2-t^{2}/2$. Then, the equations of motion (1) can be integrated from the starting point $(X_{0},Y_{0})=(x_{A},y_{A})$
@@
The found trajectory should also pass through the terminal point $(X_{T},Y_{T})=(x_{B},y_{B})$, which allows us to obtain the initial heading $\theta_{0}$ and travel time $T$. For the numerical parameters specified in Section III.2, we obtain:
@@
Figure 3 presents the optimal trajectory (white line) superimposed on the velocity field at the final time $T$. The figure also illustrate the system’s sensitivity by displaying trajectories with small errors in the initial heading $\theta_{0}$ and initial position $\bm{r}_{A}$. All trajectories remain near the optimal path and successfully approach the destination point. The complete time-dependent evolution is provided in the video, see Ref. [repository]. Based on expressions (18) and (19), it can be concluded that the initial small perturbations $\delta X_{0},\delta Y_{0},\delta\theta_{0}$ do not exhibit growth with time, indicating the neutral stability of the optimal trajectory.
Figure 3: Optimal control solution (white line) for the steady sink and time-varying vortex flow, see Sections III.2. Near-optimal trajectories are shown in cyan and blue for the initial control angle $\theta_{0}\pm\pi/360$, and in magenta and red for the initial position error $\Delta=0.025$ in $x_{A}$ and $y_{A}$.